*Present:* Si, Mullet, Tim, Polar, Odd_Bloke, Fluff_eei, Bucko, Bruce, Sadiq
* Cupboard Update
Sadiq said order was set to close on Wednesday (tomorrow), he would then calculate costs etc. and email individual signups with their personal cost.
Mullet wanted it noted that the Exec laughed at Sadiq for 10 seconds for over-complicating a calculation.
Sadiq then raised the issue of credit with suppliers. The approximate cost of the order will be around Â£250. Each member will be emailed, their order will be placed unless they specifically say they don't want it.
Tim asked why we hadn't already taken money from them.
Sadiq replied that we had no real technical mechanism yet. So we would still be relying on trust.
h4. Cupboard Update
Si told the Exec that the cupboard was OK for the Health and Safety audit. We needed to not keep any paper in there and to remove as much as possible.
On the subject of the port, Si has emailed Tech and Exec to say that we will likely only be allowed to use on computer on one port, this was a general ITS campus rule.
Si had emailed Duncan and Carol to ask for an exemption until new ports are put in but there had been no reply. Apparently every device that uses the network needs to make itself known via a unique address and there must only be one of these per port.
Mullet stated that we had no formal AUP with ITS.
Si replied that ITS have a general connection policy which was a bit uncertain in places. Everyone who uses the network is bound by ITS policies. The long and the short of the situation was that we could not use a switch, whether the policy says so or not, ITS do not want us to use one.
Mullet replied by asking how we were supposed to know these things without seeing the AUP.
Si said this was and had been available on the ITS website for ages.
*Action* Si to email ITS policy link to Exec
Exec ended the cupboard discussion - there was no real update, the Society is still waiting on ITS to audit the machines and respond to Tim's report.
h4. Tech Team
Sadiq asked what the state of play was with regards Bong's hardware. Most of it was less than 3 years old - if we had problems we might be able to get refunds or exchanges.
Mullet asked what failures there had been.
Polar said that most had come from poorly connected components which had been fixed subsequently but that there may be problems with the SATA.
Bruce said that these had been replaced before.
*Action* Polar/Tim to look at Bong, possibly run memtest to check.
h4. Gaming Update
Mullet stated that there was no update as we were waiting on the cupboard.
Si asked whether we were any further forward with investigating the technical solution to locking down the workstations.
Fluff_eei said that there had not been any real progress.
Matter is deferred until Friday XING or Tech Team meeting.
*Action* Polar to contact Brad with regards the locking down progress
Polar said that there had been an email conversation which had got confusing in places.
Polar has sent the pictures from the LAN and Accenture had replied with a form for sponsorship for next year.
*Action* Polar to fill in Accenture form.
bq=. < Si leaves for a prior engagement, meeting carries on with someone else minuting >